OpenStudy (anonymous):

the sum of 3 times a number and 2 is the same as the difference of the number and 4. Find the number. I have the answer but i dont know how to get there. -3<===== Answer

OpenStudy (lgbasallote):

let "the number" be n three times a nummber would be 3n sum of that and 2 would be 3n + 2 same as would mean = so 3n + 2 = then difference of the number and 4 would be n - 4 so the equation is 3n + 2 = n - 4 does that help?

OpenStudy (anonymous):

yes thank you very much

OpenStudy (lgbasallote):

welcome